# DBConformer: Dual-Branch Convolutional Transformer for EEG Decoding

### v1

@@ IV-F Ablation Study
IV-G Effect of Dual-Branch Modeling
To further evaluate the impact of dual-branch architecture, we conducted feature visualization experiments using $t$-SNE [47]. Features extracted by T-Conformer (temporal branch only) and DBConformer (dual-branch) were compared on four MI datasets, shown in Fig. 5. Observed that:
  • T-Conformer alone exhibits limited separability. Although T-Conformer captures meaningful temporal structures, its extracted features tend to form overlapping clusters across classes in datasets such as BNCI2014004 and Blankertz2007, which limits class discriminability.
  • DBConformer extracted more structured and well-separated features. With the integration of spatial encoding from the S-Conformer, DBConformer enhances the inter-class margins and promotes cluster compactness across all datasets, especially evident on BNCI2014001 and Blankertz2007.
  • The improvement is consistent across various MIs. The benefit of dual-branch spatio-temporal modeling is observable regardless of categories (e.g., hand vs. hand/foot), demonstrating the generalization and robustness of the proposed design.
@@ IV-G Effect of Dual-Branch Modeling
IV-H Interpretability of Channel Attention
To investigate the interpretability of the proposed channel attention module, we visualized the attention scores assigned to each EEG channel across 32 trials (a batch) from four MI datasets, illustrated in Fig. 6. We excluded BNCI2014004 from this analysis, as it only contains C3, Cz, and C4 channels and therefore lacks spatial coverage for attention comparison.
A consistent observation is that higher attention weights are repeatedly assigned to central motor-related channels, specifically C3, Cz, and C4, across all datasets. This aligns well with known neurophysiological principles of MI, where these central electrodes correspond to the sensorimotor cortex contralateral to movement imagination. Notably, even in different experimental setups, the model adaptively focuses on these physiologically relevant channels.
These results indicated that the proposed channel attention module can learn informative spatial priors in a data-driven manner, enhancing both performance and interpretability.
Figure 6: Attention scores of each EEG channel calculated by the proposed channel attention module on (a) BNCI2014001, (b) Zhou2016, (c) BNCI2014002, and (d) Blankertz2007 datasets. The sum of scores from all channels of each sample is 1.
TABLE VII: Comparison of model complexity and performance metrics across DBConformer and nine baseline models on the BNCI2014001 dataset.
Metrics CNN CNN-Mamba CNN-Transformer
EEGNet SCNN DCNN FBCNet ADFCNN IFNet SlimSeiz CTNet EEG Conformer DBConformer (Ours)
# Model Parameters 1,406 46,084 320,479 7,042 4,322 7,748 27,650 27,284 789,668 92,066
Training Duration (s) 73.64 76.44 90.97 70.45 97.84 61.89 88.89 105.95 223.36 203.79
Accuracy (%) 73.64 72.22 73.21 72.56 71.76 74.52 69.62 73.40 73.07 77.67
IV-I Model Complexity and Performance Analysis
@@
Table VII presents a comparative analysis of trainable parameter counts, training duration, and classification accuracy across DBConformer and nine baseline models. Observed that:
@@
  • Compared to lightweight CNN models, DBConformer introduces additional complexity due to the dual branches. Nonetheless, the trade-off in model size and training time is well-compensated by superior performance.
  • DBConformer achieves the highest classification accuracy, surpassing all competing CNN and Conformer models. Despite incorporating a dual-branch Conformer design, it maintains a moderate number of parameters, with over 8$\times$ fewer parameters than the high-capacity EEG Conformer.
  • Overall, DBConformer demonstrates an effective balance between model capacity, computational efficiency, and accuracy, highlighting its suitability for EEG decoding.
@@ V Conclusion
This paper proposed DBConformer, a dual-branch convolutional Transformer network tailored for EEG decoding. It integrates a T-Conformer to model long-range temporal dependencies and an S-Conformer to enhance inter-channel interactions, capturing both temporal dynamics and spatial patterns in EEG data. A channel attention module further refines spatial characteristics by assigning data-driven importance to EEG channels. Comprehensive experiments on five MI datasets and two seizure detection datasets under three settings demonstrate that DBConformer consistently outperforms 10 competitive baseline models, with over 8$\times$ fewer parameters than the high-capacity EEG Conformer baseline. Further, the visualization results confirm that the features extracted by DBConformer are physiologically interpretable and aligned with sensorimotor priors in MI.
In future work, DBConformer can be extended to support multi-view information fusion and plug-and-play real-time BCIs. Its superior performance and interpretability make it reliable for robust and explainable EEG decoding.

### v2

@@ IV-G Ablation Study
IV-H Effect of Dual-Branch Modeling
To further evaluate the impact of dual-branch architecture, we conducted feature visualization experiments using $t$-SNE [65]. Features extracted by T-Conformer (temporal branch only) and DBConformer (dual-branch) were compared on four MI datasets, shown in Fig. 5. Observe that:
  • T-Conformer alone exhibited limited separability. Although T-Conformer captured meaningful temporal structures, its extracted features tended to form overlapping clusters across classes in datasets such as BNCI2014004 and Blankertz2007, which limited class discriminability.
  • DBConformer extracted more structured and well-separated features. With the integration of spatial encoding from the S-Conformer, DBConformer enhanced the inter-class margins and promotes cluster compactness across all datasets.
  • The improvement was consistent across various types of MI classification tasks. The benefit of dual-branch spatio-temporal modeling was observable regardless of categories (e.g., hand vs. hand/foot), demonstrating the generalization and robustness of the proposed design.
@@ IV-H Effect of Dual-Branch Modeling
IV-I Visualization of Spatio-Temporal Self-Attention
To further examine the interpretability of DBConformer, Fig. 6 visualized the self-attention matrices learned in both temporal and spatial branches on BNCI2014001, BNCI2014002, and OpenBMI datasets. Observe that:
  • The temporal self-attention maps in Fig. 6(a) exhibited diagonal activations, which corresponded to the model’s focus on local neighboring time windows, as well as noticeable off-diagonal activations that captured long-range temporal dependencies. Across three datasets, DBConformer emphasized short-term EEG dynamics and integrated global temporal information, enabling a more comprehensive representation of MI.
  • The spatial self-attention maps in Fig. 6(b) revealed both diagonal activations corresponding to self-focus on individual channels, and distinct off-diagonal activations reflecting inter-channel interactions. In three datasets, channels over the sensorimotor cortex (e.g., C3, Cz, C4) exhibited the top attention intensities, indicating that DBConformer effectively captured physiologically meaningful intra-channel and inter-channel dependencies.
  • Furthermore, channel importance based on incoming off-diagonal attention weights is quantified in Fig. 6(c). Motor-related channels consistently dominated the importance ranking, confirming that DBConformer robustly identifies physiologically relevant channels.
Figure 6: Visualizations of temporal and spatial self-attention across three MI datasets. (a) Temporal attention heatmap of time windows; (b) Spatial attention heatmap of EEG channels; (c) Channel importance of electrodes.
IV-J Model Complexity and Performance Analysis
@@
Table VIII presents a comparative analysis of trainable parameter counts, training duration, and classification accuracy across DBConformer and nine baseline models. Observe that:
@@
  • Compared to lightweight CNN models, DBConformer introduced additional complexity due to the dual branches. Nonetheless, the trade-off in model size and training time was well-compensated by superior performance.
[…]
  • IFNet and EEGNet achieved the shortest inference times among CNN-based models. Importantly, DBConformer achieved the fastest inference speed among CNN-Transformer hybrids, requiring only 8.7 ms per batch, which meets the real-time requirements for online BCI applications.
TABLE VIII: Comparison of model complexity and performance metrics across DBConformer and twelve baseline models on the BNCI2014001 dataset.
Metric CNN CNN-Mamba CNN-Transformer
EEGNet SCNN DCNN FBCNet ADFCNN IFNet SlimSeiz CTNet EEG Confor. MSCFor. TMSA-Net MSVTNet DBConformer
# Model Parameters 1,406 46,084 320,479 7,042 4,322 7,748 27,650 27,284 789,668 150,626 17,869 72,892 92,066
Training Duration (s) 73.64 76.44 90.97 70.45 97.84 61.89 88.89 105.95 223.36 151.63 96.33 159.80 203.79
Inference Latency (s) 4.5 $e^{-3}$ 1.3 $e^{-2}$ 4.8 $e^{-3}$ 1.3 $e^{-2}$ 7.8 $e^{-3}$ 4.2 $e^{-3}$ 1.7 $e^{-2}$ 1.3 $e^{-2}$ 1.9 $e^{-2}$ 1.6 $e^{-2}$ 8.8 $e^{-3}$ 9.0 $e^{-3}$ 8.7 $e^{-3}$
Accuracy (%) 73.64 72.22 73.21 72.56 71.76 74.52 69.62 73.40 73.07 75.00 76.67 74.60 77.67
IV-K Sensitivity Analysis on Architectural Design
We further conducted a sensitivity analysis to explore how architectural design affects the DBConformer performance.
IV-K1 Effect of Depth of T-Conformer and S-Conformer
We varied the number of Transformer encoder layers in both T-Conformer and S-Conformer from 1 to 8, and evaluated the classification accuracy on BNCI2014001 and Zhou2016 datasets. Results in Fig. 7(a) showed that performance does not monotonically increase with depth. Instead, both datasets achieve the best performance with two Transformer layers. When the depth exceeds 4 layers, accuracy tends to decrease slightly, suggesting that excessive stacking of attention layers may lead to overfitting on the limited training samples in MI datasets. In contrast, shallow configurations achieve a favorable trade-off between model complexity and generalization.
IV-K2 Effect of the Number of Attention Heads
Constrained by the embedding size of $D=40$, which requires the number of heads to be a divisor of $40$, we evaluated the number of attention heads 1,2,4,5,8,10{1,2,4,5,8,10} on BNCI2014001 and Zhou2016 datasets. As shown in Fig. 7(b), the results remain relatively stable across different settings, with DBConformer consistently outperforming baseline models. Notably, $2$ heads achieved the highest accuracy on both datasets. Increasing the number of heads beyond $4$ did not yield further improvements and sometimes led to a slight decrease in performance, suggesting that excessive partitioning of the embedding dimension may dilute feature representations.
Figure 7: Sensitivity analysis on architectural design: (a) Number of Transformer layers in T-Conformer and S-Conformer, (b) Number of attention heads in T-Conformer and S-Conformer. Each bar represents the mean accuracy across subjects, with error bars denoting the standard deviation.
@@ V Conclusion
This paper proposed DBConformer, a dual-branch convolutional Transformer network tailored for EEG decoding. It integrates a T-Conformer to model long-range temporal dependencies and an S-Conformer to enhance inter-channel interactions, capturing both temporal dynamics and spatial patterns in EEG data. Comprehensive experiments confirmed its effectiveness. In future work, DBConformer can be extended to support multi-view information fusion and plug-and-play real-time BCIs. Its superior performance and interpretability make it reliable for robust and explainable EEG decoding.